[figure: 9.7 cm]

HEAVY-FERMION BEHAVIOUR OF LiV_{2}O_{4}

R.J. Radwański^{a,b}, Z. Ropka^{b} and R. Michalski^{a,b} aInst. of Physics, Pedagogical University, 30-084 Kraków, bCenter for Solid State Physics, św. Filip 5, 31-150 Kraków Poland.

LiV_{2}O_{4} has been announced as the first heavy-fermion 3d-ion oxide (Phys.Rev.Lett. 78 (1997) 3729). Its Sommerfeld coefficient γ of 0.42 J/K^{2}mol is surely not very large (CeAl_{3}: 1600 mJ/K^{2}mol but UPt_{3}: 430 mJ/K^{2}mol) but it is exceptionally large for a 3d-ion system. It shows no magnetic order down to 0.02 K. Though χ exhibits the Curie-Weiss law above 100 K with S≃1/2, it stabilizes at low temperatures. We have studied properties of the V^{3+} and V^{4+} ions in LiV_{2}O_{4}. We have completed evidence that these heavy-fermion (h-f) like properties can result from particularities of the fine electronic structure of the d^{1} system and the breaking of the time reversal symmetry in the atomic scale. Finally we would like to point out that the occurence of the h-f behaviour for a 3d-ion system, with an odd number of d electrons, has been predicted by the crystal-field model. This model has been originally developed for the f-electron systems and the finding of the h-f state in LiV_{2}O_{4} we are taking as the nice confirmation of the model. Finally, we would like to mention that according to our results the excitations involved in the h-f state in LiV_{2}O_{4} are of the spin nature in contrary to often discussed chargefluctuation excitations.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : R.J. Radwanski

Address for correspondence : Center for Solid State Physics, św. Filip 5, 31-150 Kraków

Email address : sfradwan@cyf-kr.edu.pl